OpenStudy (anonymous):

@BlackLabel What Is An Equation Of A Parabola With The Given Vertex And Focus? Vertex: (-2,5); Focus: (-2,6)

OpenStudy (anonymous):

wtf is a focus??? Lemme google that

OpenStudy (anonymous):

Lmao, Alright

OpenStudy (anonymous):

This is new for me. I have never come across it Using the focus and the vertex we use the following formula to find our equation (x – h)2 = 4p(y – k) h= x-value of our vertex k= y-value of the vertex p= the distance btwn the focus and the vertex \( \to y_1-y_2\)

OpenStudy (anonymous):

What is our h,k and p?

OpenStudy (anonymous):

h= -2 k= 5 p= 1?

OpenStudy (anonymous):

or p=-1?

OpenStudy (anonymous):

p= focus -Vertex ----> 6-5=1

OpenStudy (anonymous):

Ok so it is p=1

OpenStudy (anonymous):

\[(x – h)^2 = 4p(y – k) \] Plugging in the values we get \[(x+2)^2=4(1)(y-5)\] \[(x+2)^2=4(y-5)\] Divide both sides by 4 \[\frac{1}{4}(x+2)^2=(y-5)\] Add 5 to both sides \[\frac{1}{4}(x+2)^2+5=y\] And there we go

OpenStudy (anonymous):

I Love You ♥

OpenStudy (anonymous):

Love you too ♥

OpenStudy (anonymous):

Can You Help Me With One More? C:

OpenStudy (anonymous):

Just Juan

OpenStudy (anonymous):

ya sure go for it

OpenStudy (anonymous):

A Radio Station Has A Broadcast Area In The Shape Of A Circle With Equation x^2+y^2=5,625, Where The Constant Represents Square Miles. a. Find The Intercepts Of The Graph And State The Radius In Miles. b. What Is The Area Of The Region In Which The Broadcast From The Station Can Be Picked Up?

OpenStudy (anonymous):

Ok x-intercept is when y=0 so lets plug y=0 into the equation and see what x equals \[x^2+(0)^2=5625\] Square root both sides \[x=\sqrt{5625}=75\] So the x intercept is (75,0)

OpenStudy (anonymous):

Since it is a circle we can assume that the y intercept is 75 too but lets do it algebraically anyways The y-intercept is when x=0 So lets plug x=0 into the equation and see what y equals \[(0)^2+y^2=5625\] Square root both sides \[y=\sqrt{5625}=75\] So the y-intercept is (0,75)

OpenStudy (anonymous):

Ohhh fu.ckkkkk hmmmm gotta add smth

OpenStudy (anonymous):

Lmao

OpenStudy (anonymous):